OpenStudy (anonymous):

a(3tx-2b)=c(dx-2) find x

OpenStudy (turingtest):

distribute a and c what do you get?

OpenStudy (anonymous):

3txc-2bc=dcx-2c?

OpenStudy (turingtest):

how did you get a c on the left?

OpenStudy (turingtest):

you have an a on the left and a c on the right...

OpenStudy (anonymous):

didn't meant to press that........ 3tx-2bc=dcx-2c <--- meant that

OpenStudy (turingtest):

there is still no a on the left, and there is still a c just simplify the left: a(3tx-2b)=?

OpenStudy (anonymous):

3txa-2ab

Directrix (directrix):

3txa-2ab = c(dx-2) What does c(dx-2) equal when the "c" is distributed over the (dx-2)?

OpenStudy (anonymous):

dxc-2c

Directrix (directrix):

3txa-2ab = c(dx-2) 3txa -2ab = cdx -2c Get the terms containing x together on one side of the equation and the terms not containing x on the other side. 3txa - cdx = 2ab - 2c (3ta)x - (cd) x = 2ab - 2c Check that. Then, factor x from the left side to get: x [ 3at - cd ] = 2ab - 2c Then, what do you think should happen next?

OpenStudy (anonymous):

divide both sides by 3at-cd so x is left alone